# Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach

### v2

@@ 3.2 Dynamic Stackelberg Mean Field Games
where the followers’ state-action distribution $L^{f}$ satisfies the following Mckean-Vlasov equation:
$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l}).$ (3)
In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic Stackelberg mean field game. This modeling approach aims to learn the optimal policy of the leader, namely the social planner, while considering the followers’ best response over discrete timesteps (Definition 3.2).
3.3 Metrics for Dynamic SMFGs
We will utilize the following three metrics to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the followers’ policy from the best response to the leader’s policy; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
Leader’s Payoff
Exploitability
Exploitability is a widely used metric for evaluating convergence and measuring the distance between the representative follower’s policy $\pi^{f}$ and the best response ${\pi^{f}}^{*}$ for the leader’s policy $\pi^{l}$. Formally, it quantifies the extent to which a single follower can improve their payoff by deviating from the population’s behavior to adopt a different policy. Given the leader’s policy $\pi^{l}$ and the followers’ state-action distribution $L^{f}$, the exploitability $\mathcal{E}(\pi^{f};\pi^{l})$ of the follower’s policy $\pi^{f}$ is defined by comparing the value functions induced by $\pi^{f}$ and its best response $\pi^{f*}$:
$\mathcal{E}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L^{f})-J^{f}(\pi^{l},{\pi^{f}},L^{f}).$
Social Optimum and Social Welfare
In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 40]. Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions of the $N$ followers in Equation 1 when $N\to\infty$:
$\mathcal{SW}(\pi^{l},\pi^{f})=\sum^{N}J^{f}\left(\pi^{l},\pi^{f},L^{f}\right).$
@@ 4 Stackelberg Mean Field Reinforcement Learning
However, solving for the leader’s optimal policy in dynamic SMFGs, as defined in Definition 3.2, presents several challenges, such as vast state and action spaces, strategic dependencies and information asymmetry between the leader and followers, and potentially conflicting objectives including the leader’s payoff, followers’ utilities, and social welfare. These complexities make devising a model-based algorithm with theoretical proof extremely difficult. Previous works have attempted simplifications using the myopic followers setting [63] or linear-quadratic MFGs [6], but this approach fails to align with macroeconomic policy-making issues. Consequently, we have designed the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to address the above limitations. This is a model-free approach that operates independently of environmental transition information.
@@
Considering the dynamic SMFGs with a leader and $N$ homogeneous follower agents, at any time $t\in\{0,\ldots,T-1\}$, the joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next state $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$. The experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording the experiences of all agents. Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$). The SMFRL algorithm incorporates several settings (shown in Figure 2) to address the aforementioned challenges:
@@
To mitigate the problem of non-stationarity, we consider a central critic Q function $Q(\mathbf{s}_{t},\mathbf{a}_{t}):\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}\times\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}\to\mathbb{R}$ for each agent. However, due to the high dimensionality of the joint state and action, the standard Q network is infeasible to learn. Thus, the SMFRL algorithm employs a mean field approximation, characterizing the followers’ state-action pairs $(\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t})$ by using the followers’ population distribution $L^{f}_{t}$, which can be presented through various metrics such as mean, variance, probability density functions, skewness, and kurtosis. The theoretical proof is detailed in Appendix F. This approach yields $Q^{l}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ and $Q^{f}(s^{f}_{t},a^{l}_{t},a^{f}_{t},L^{f}_{t})$ for evaluating the leader and follower agents respectively. The mean field approximation effectively reduces the dimensionality of the inputs to the Q function, thereby facilitating the learning of the critic Q network and ensuring the Q function can evaluate leader and follower policies based on the overall follower distribution.
@@
In SMFRL, the leader’s policy $\pi_{\theta^{l}}$, is optimized using a deterministic policy gradient approach [54]. The policy gradient for the leader can be estimated as:
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right],$
where ${Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$, the leader’s action-value function, is approximated by a neural network parameterized by $\phi^{l}$. This network is periodically updated to minimize the loss:
$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
where the target value $y^{l}_{t}$ is given by:
$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$
and $\phi^{l}_{-}$ represents the parameters of the target Q network, $L^{f}_{t+1}$ is determined based on Mckean-Vlasov equation (3), and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right].$
Followers’ Update for Best Response
To enable follower agents to effectively learn the best response policy $\pi^{f*}$ to the leader’s policy $\pi^{l}$, we employ multiple updates for followers’ shared policy and Q networks in the SMFRL algorithm. Each update cycle refines the policy network $\pi_{\theta^{f}}$ for the followers using the policy gradient method:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f}_{t})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
where the action-value function ${Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$, parameterized by $\phi^{f}$, is periodically updated by minimizing the mean squared error (MSE) loss:
$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left[\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$ (4)
$y^{f}_{t}=$ $r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
with the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, and the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$. The joint actions of followers $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},\ldots,a^{fN}_{t+1}\}$ are determined with each action $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
The pseudocode for the SMFRL algorithm (1) is presented in Appendix A.
Table 1: Evaluation of SMFRL and five baselines on 3 metrics for dynamic Stackelberg mean field games with different followers’ numbers.
Metrics Followers’ Optimal SMFRL MF-MARL MADDPG IDDPG DDPG U.S. Policy
Number Value +Real +Real
Leader’s Payoff N=100 \ 3387.36 -76.82 -282.62 -701.61 2840.07 1718.08
N=1000 \ 3411.16 -668.03 121.18 -767.53 2874.19 -796.4
Exploitability N=100 0. 0.000107 0.002190 0.007908 0.001333 0.009885 0.032490
N=1000 0. 0.000158 0.003481 0.035613 0.000565 0.003002 0.010522
Social Welfare N=100 100 96.87 82.02 55.74 83.09 94.5 79.23
N=1000 1000 968.94 834.93 645.38 874.53 943.82 334.79
@@ 5 Experiment
In this section, we conduct experiments within the TaxAI environment to answer the following three questions: (1) How effective is our SMFRL algorithm in solving dynamic SMFGs? (2) For macroeconomic policy-making issues, how does our method compare to real-world strategies, economic models, and existing AI-based methods? (3) Given the likely real-world scenario where some households do not adopt the SMFG household policy learned by our method, does our approach remain effective?
@@
TaxAI [41] is one of the rare reinforcement learning environments designed to study optimal macroeconomic policies. It simulates dynamic interactions between large-scale households and the government, making it particularly well-suited for experimenting with SMFGs. TaxAI models the optimal tax policy problems where the government optimizes tax policies to promote economic growth (i.e. GDP) and social equity (i.e. Gini). In response, households optimize their labor supply and savings strategies to maximize their long-term utility.
5.1 Solving Dynamic SMFGs
Metrics and Baselines
We compare several classic multi-agent reinforcement learning (MARL) algorithms based on three metrics: leader’s payoff, exploitability, and social welfare, as defined in Section 3.3. Optimal Value: The ideal value of exploitability is 00, indicating that the follower’s policy achieves the best response to the leader’s policy. The upper bound of Social Welfare is $N$, given that household utility in TaxAI cannot exceed 1. For Leader’s Payoff, no maximum has been established. MF-MARL [61]: Utilizes the DDPG [36] algorithm for the leader and a mean-field MARL approach for the followers. MADDPG [38] and IDDPG [17]: Both feature shared actor and critic networks among followers. DDPG+Real: DDPG for the leader and behavior cloning [57] from real data for followers. U.S. policy+Real: U.S. federal tax policy for the leader with behavior cloning for the followers.
Results Analysis
Table 1 presents the performance metrics for SMFRL and five baselines, alongside their optimal values, averaged over 10 evaluations of models trained with at least five seeds. SMFRL notably outperforms the baselines in achieving higher leader’s payoff, demonstrating its effectiveness in optimizing the leader’s objective. The DDPG+Real method, using a real-data trained policy, attains a substantial payoff but exhibits a significant gap from the households’ optimal response, as indicated by its higher exploitability. In contrast, SMFRL significantly reduces exploitability for both $N=100$ and $N=1000$, suggesting that its policy for households closely approximates the best response to the leader’s strategy, which is crucial for solving Stackelberg mean field games. Moreover, SMFRL closely approaches the optimal social welfare value, indicating that its solutions are the most socially optimal among the considered baselines. In summary, the SMFRL algorithm effectively learns policies that maximize the leader’s objective while ensuring that followers attain the best response to the leader’s policy, closely approaching the social optimum. The training curves are presented in Figure 4, and the test results for macroeconomic indicators are detailed in Table 1 in Appendix D.
5.2 Macroeconomic Policy-Making
Table 2: Test results 6 tax policies across five macroeconomic indicators at two numbers of followers, $N=100$ and $N=1000$.
Households’ Policies Macroeconomic Indicators
@@ 5.2 Macroeconomic Policy-Making
Number Per Capita GDP Social Welfare Income Gini Wealth Gini Years
$\mathbf{N=100}$ Free Market 1.37e+05 32.97 0.89 0.92 1.10
Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55
AI Economist 1.26e+05 72.81 0.88 0.91 1.00
SMFG 9.59e+12 96.87 0.52 0.51 300.0
$\mathbf{N=1000}$ Free Market 1.41e+05 334.79 0.90 0.93 1.00
Saez Tax 6.35e+11 498.88 0.68 0.73 100.578
US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
AI Economist-BC N/A N/A N/A N/A N/A
AI Economist N/A N/A N/A N/A N/A
SMFG 1.10e+13 968.94 0.54 0.53 300.00
Metrics and Baselines
@@
In addressing macroeconomic policymaking, we utilize classical macroeconomic indicators from the real world to evaluate the performance of different policies. Per Capita GDP indicates the level of economic development, while Income Gini and Wealth Gini measure the inequality in households’ income and wealth, respectively; a lower Gini index suggests greater social equality. The Years metric represents the sustainable duration of an economy, capped at 300 years; economies may terminate earlier due to excessive inequality or inadequate productivity to meet the populace’s needs. We compare several methods for solving macroeconomic policies: the Free Market [5] policy entails no government intervention; the Saez Tax [50] represents a well-known economic approach, often advocated for specific tax reforms in practice. The U.S. Federal Tax is the real policy from the OECD data  [^2] https://data.oecd.org/. In these approaches, households employ strategies based on behavior cloning from real data. Additionally, the AI Economist [62] represents an AI-based policy using a two-level Proximal Policy Optimization (PPO) framework. The AI Economist-BC method leverages behavior cloning policy on real-world data for pretraining.
Results Analysis
Table 2 presents the average values of macroeconomic indicators across multiple test rounds for different algorithms under varying numbers of households. The results reveal that the Saez Tax method achieves the lowest income Gini and wealth Gini when the number of households is small ($N=100$). However, our SMFG method yields higher per capita GDP and social welfare. At $N=1000$, the advantages of the SMFG algorithm become even more pronounced compared to other methods. It secures the longest duration (Years), the highest GDP, and the lowest income Gini and wealth Gini. The social welfare derived from the SMFG approach closely approaches the optimal value of 1000. This demonstrates that our method can learn the macroeconomic policy that simultaneously optimizes economic growth, social equity, and social welfare, surpassing existing economic approaches like the Saez Tax and AI-based policies (i.e. AI Economist). We provide the corresponding training curves in Figure 5 in Appendix D.
5.3 Effectiveness Analysis in Diverse Households

### v3

@@ 3 Stackelberg Mean Field Game Modeling
where the followers’ state-action distribution $L_{t}$ satisfies the following McKean-Vlasov equation:
$L_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})$
$P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L_{t})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l},L_{t}).$
In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic SMFG. Under this modeling framework, we can optimze the macroeconomic policy of the leader, namely the government, while considering the followers’ best response over discrete timesteps (Definition 2).
@@ 4 Stackelberg Mean Field Reinforcement Learning
To solve macroeconomic policymaking under Dynamic SMFG framework, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which is a centralized training with decentralized execution (CTDE) algorithm. We introduce the leader’s policy $\pi^{l}(s^{l})$ and a central critic Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for the leader agent, and the shared policy $\pi^{f}(s^{f},a^{l})$ and $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ for each follower agent. SMFRL algorithm comprises two key components: mean field approximation and leader-follower update (see Figure 2).
@@
Due to the dimensions of joint state $\mathbf{s}^{f}$ and action $\mathbf{a}^{f}$ scaling with the number of follower agents, the standard Q-function $Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ becomes infeasible to learn when dealing with a large number of followers. Therefore, we factorize the standard Q-function using pairwise local interactions between the leader and each follower (Yang et al. 2018):
$Q^{l}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ $=\frac{1}{N}\sum_{i}^{N}\tilde{Q}^{l}(s^{l},a^{l},s^{fi},a^{fi})$ (2)
$\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{l}(s^{l},a^{l},s^{f},a^{f})$
Similarly, the $i$-th follower’s Q-function $Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})$ can be factorized using local interactions between the leader, the $i$-th follower, and each of the remaining followers:
$Q^{f}(s^{l},a^{l},\mathbf{s}^{f},\mathbf{a}^{f})\approx\mathbb{E}_{(s^{f},a^{f})\sim L}\tilde{Q}^{f}(s^{l},a^{l},s^{fi},a^{fi},s^{f},a^{f})$
where $(s^{fi},a^{fi})$ and $(s^{f},a^{f})$ are the state-action pair for the $i$-th follower and another follower agent, respectively. We use $\hat{L}(\pi^{f},a^{l})$, a finite set of state-action pairs sampled under the follower’s policy $\pi^{f}$ and the leader’s action $a^{l}$, to construct an empirical distribution that approximates the state-action distribution $L(s^{f},a^{f})$.
$\hat{L}(\pi^{f},a^{l})=\left\{(s^{fi},a^{fi})\right\}_{i=1}^{N},\text{ where }a^{fi}=\pi^{f}(s^{fi},a^{l}).$
Leader-follower Update
To enhance the convergence and stability, we propose the leader-follower update: first, by fixing the leader’s policy, we train the followers’ shared policy and Q-networks towards the best response; subsequently, based on the followers’ policies, we optimize the leader’s policy and Q net, alternating these steps until convergence. We measure the distance between the agents’ policies and their best responses by using exploitability.
Based on mean field approximation, we will train these networks $\pi_{\theta^{l}}$ and $\tilde{Q}_{\phi^{l}}$ for the leader agent, shared $\pi_{\theta^{f}}$ and $\tilde{Q}_{\phi^{f}}$ for the follower agents, with parameters $\theta^{l}$, $\phi^{l}$, $\theta^{f}$, and $\phi^{f}$. To ensure training stability, we introduce target networks with parameters $\theta^{l}_{-}$, $\phi^{l}_{-}$, $\theta^{f}_{-}$, and $\phi^{f}_{-}$. At any step $t\in\{0,\ldots,T\}$, the tuple $(s^{l}_{t},a^{l}_{t},\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t},s^{l}_{t+1},\mathbf{s}^{f}_{t+1},r^{l}_{t},\mathbf{r}^{f}_{t})$ is stored in the replay buffer $\mathcal{D}$ for training. When the number of followers is particularly large, a finite subset of state-action pairs $\{(s^{fi}_{t},a^{fi}_{t})\}_{i=1}^{N}$ can be selected for storage.
Followers’ Update for Best Response
Given leader’s policy $\pi_{\theta^{l}}$, the followers’ policy network $\pi_{\theta^{f}}$ is updated using a deterministic policy gradient approach (Silver et al. 2014). The followers’ policy gradient is estimated:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
$\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})\nabla_{a^{fi}_{-}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{-},s^{f}_{t},a^{f}_{t})\Big{]}$
where $a^{fi}_{-}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})$, $\hat{L}_{\mathcal{D}}$ denotes state-action pairs sampled from replay buffer $\mathcal{D}$. The action-value function $\tilde{Q}_{\phi^{f}}$ is updated by minimizing the mean squared error loss:
$\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
$y^{fi}_{t}=r^{fi}_{t}+$
$\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}_{-}},a^{l}_{t+1})}\tilde{Q}_{\phi^{f}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{fi}_{t+1},a^{fi}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
where $a^{l}_{t+1}=\pi_{\theta^{l}}(s^{l}_{t+1})$, $a^{fi}_{t+1}=\pi_{\theta^{f}_{-}}(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}$ $\Big{[}\left(y^{fi}_{t}-\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\right)$
$\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{l}_{t},a^{l}_{t},s^{fi}_{t},a^{fi}_{t},s^{f}_{t},a^{f}_{t})\Big{]}.$
@@
Given followers’ policy $\pi_{\theta^{f}}$, the leader’s policy $\pi_{\theta^{l}}$ is optimized by DPG approach, and the leader’s policy gradient is estimated:
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D},(s^{f},a^{f})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{-})}$
$\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l},a^{l}_{-},s^{f},a^{f})\right],\text{where }{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}.$
This network is periodically updated to minimize the loss:
$\mathcal{L}(\phi^{l})=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D}\\ (s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}\end{subarray}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)^{2}\right]$
The target value $y^{l}_{t}$ is given by:
$y^{l}_{t}=r_{t}^{l}+\gamma\mathbb{E}_{(s^{f}_{t+1},a^{f}_{t+1})\sim\hat{L}(\pi_{\theta^{f}},a^{l}_{t+1})}\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},s^{f}_{t+1},a^{f}_{t+1})$
where $a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})$, and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}$ $\mathcal{L}(\phi^{l})=\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},\mathbf{s}_{t+1}\sim\mathcal{D},(s^{f}_{t},a^{f}_{t})\sim\hat{L}_{\mathcal{D}}}$
$\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},s^{f}_{t},a^{f}_{t})\right].$
The pseudocode for the SMFRL algorithm 1 is presented in Appendix A.
@@ 5 Experiment
In this section, we conduct experiments of the SMFG method, which is modeled by dynamic SMFGs and solved by the SMFRL algorithm, within the TaxAI (Mi et al. 2023) environment to answer the following questions:
1. For macroeconomic policy-making issues, how does our SMFG method compare to real-world policies, classic economic methods, and existing AI-based policy?
2. Are both the leader-follower update and mean-field approximation modules of the SMFG method essential?
3. Given the likely real-world scenario where some households do not adopt the SMFG policy, does the SMFG approach remain effective?
@@
TaxAI is a simulation environment based on real data designed to study optimal taxation policies. It simulates the dynamic interactions between the government and large-scale households, to promote economic growth (i.e., GDP) as the objective.
@@
Tax Policies Per Capita GDP Social Welfare Income Gini Wealth Gini Years
$100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$ $100$ $1000$
Free Market 1.37e+05 1.41e+05 32.97 334.79 0.89 0.90 0.92 0.93 1.10 1.00
Saez Tax 2.34e+12 6.35e+11 73.82 498.88 0.21 0.68 0.38 0.73 300.00 100.58
US Federal Tax 4.88e+11 1.41e+05 94.19 351.17 0.40 0.89 0.40 0.93 289.55 1.00
AI Economist-BC 4.24e+12 N/A 97.24 N/A 0.54 N/A 0.52 N/A 299.55 N/A
AI Economist 1.26e+05 N/A 72.81 N/A 0.88 N/A 0.91 N/A 1.00 N/A
SMFG 9.59e+12 1.10e+13 96.87 968.94 0.52 0.54 0.51 0.53 300.00 300.00
Table 1: Assessment results of 6 macroeconomic policies for different numbers of households N=100 and N=1000.
Figure 3: Dynamic response curves of 3 macroeconomic policies to economic shocks at step 100. The SMFG policy (dark blue line) exhibits the least fluctuation and the fastest recovery in key indicators, indicating superior dynamic response capabilities.
Figure 4: Ablation studies of SMFG method based on macroeconomic indicators for households’ numbers N=100 and N=1000.
Comparison of Macroeconomic Policies
This section compares the tax policy derived from SMFG with existing taxation policies in two aspects: (1) Macroeconomic Policy Evaluation: We assess the performance of different policies using key macroeconomic indicators. (2) Dynamic Response: At step 100, we simulate an economic shock caused by a financial crisis to evaluate the dynamic response capabilities of the different tax policies.
Baselines and Metrics
@@
We select four different types of tax policy baselines, including the Free Market (Backhouse 2005) policy, which entails no government intervention; the Saez Tax (Saez 2001), a well-known economic approach often advocated for specific tax reforms in practice; the U.S. Federal Tax, which is the actual policy derived from OECD data https://data.oecd.org/; and the AI Economist (Zheng et al. ), a popular AI-based policy utilizing a two-level Proximal Policy Optimization (PPO) framework. Additionally, the AI Economist-BC method employs behavior cloning on real-world data for pretraining. In the TaxAI environment, we evaluate the performance of these policies using key macroeconomic indicators derived from real-world data. Per Capita GDP reflects the level of economic development, while Income Gini and Wealth Gini measure inequality in household income and wealth, respectively—a lower Gini index indicates greater social equality. The Years metric represents the sustainable duration of an economy, with a maximum cap of 300 years. Average Wealth, Income, and Consumption are crucial assessment metrics related to financial crises.
Macroeconomic Policy Evaluation
In the context of optimizing economic growth, our SMFG method achieves the highest per capita GDP, social welfare, and the longest duration of outcomes across both $N=100$ and $N=1000$ scenarios. Table 1 shows the average values of macroeconomic indicators across 10 test runs for different algorithms under varying household numbers. Although the Gini index for SMFG is higher than that of the Saez Tax when $N=100$, this is acceptable when optimizing GDP, as SMFG’s per capita GDP is nearly 4 times that of the Saez Tax. When $N=1000$, SMFG outperforms all other metrics, indicating that the advantages of our SMFG method increase with the number of households. Figure 7 in the Appendix provides the corresponding training curves.
Dynamic Response
We simulate an economic shock similar to a financial crisis in the TaxAI environment: at step 100, all households lose 50% of their wealth. Due to the short duration (Years) under the Free Market and AI economist strategies, we focus on comparing the Saez Tax, US Federal Tax, and SMFG policy.
From Figure 3, we find that the SMFG policy (dark blue line) demonstrates the fastest dynamic response across all indicators: Social Welfare remains largely unaffected, while Average Income and Average Consumption recover to their original levels within about 5 steps. Wealth Gini and Average Wealth return to their previous levels within about 15 steps. This level of recovery is unmatched by any other tax policy.